# On an entropy inequality for quadratic forms and applications

### v1

@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
Indeed, our proof of Theorem 1.1 works for any real quadratic form $\phi$ that is congruent over $\mathbb{Q}$ to a real diagonal form, i.e., $\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\Xi=\begin{bmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\end{bmatrix}\in GL_{2}(\mathbb{Q})$, together with the constant $\epsilon$ depending both on $s$ and $\Xi$.
@@
Let $X$, $Y$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$. Let $(X^{\prime},Y^{\prime})$ be i.i.d. with $(X,Y)$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
@@
To compare Theorems 2.32 and 6.1, one may take $W=(X,Y)$ in Theorem 2.32. The random variables $X$ and $Y$ are required to be independent in Theorem 6.1, but may be dependent in Theorem 2.32.
@@
[…]
@@
Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $20<s\leq 1/$ 2 and $C\geq 1$. Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$. Then for any $\zeta>0$, we have that
@@
Let $10<s<$ 1 and $C\geq 1$. Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$. For any $30<\gamma<1/$ 3, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
@@
By construction, $\text{dist}(U_{1},U_{2})\geq 2^{-k}\gg(\gamma/C)^{1/s}$. And by Lemma B.2, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman. This completes the proof. ∎
@@
Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later. We apply Lemma 6.3 on the distribution measure of $X$, denoted by $\mu$. Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $21/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/$ 2.
@@
For simplicity, write $a=\mu(U_{1})$ and $b=\mu(U_{2})$, with $21/2-\gamma\leq a,b\leq 1/$ 2. It follows that
@@ 6.2.1 Material from Shmerkin and Wang
[…]
@@
$\nu(Y_{x}\cap T)\leq K\cdot r^{t}\text{ for all }r>\delta\text{ and all open $r$-tubes $T$ through $x$}.$
@@
[Variant of Proposition 5.11 in [20]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in $[0,1],$then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
@@
[…]
@@
[…]
@@
Next we introduce a particularly useful a relaxation of the notion of an $(s,C)$-Frostman measure. In a metric space $X,$ we will call a measure $\mu\in\mathcal{P}(X)$ a $(\delta,s,C)$-measure if $\mu(B_{r})\leq C\cdot r^{s}$ for all $r\in[\delta,1].$ Given $t\in[0,d]$, $s\in[0,k],$ a function $D:[0,d]\to[0,k],$ and $C\geq 1$, we say that a measure $\rho\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$ is $(D;C)$-adapted at scale $\delta$ if whenever $\epsilon>0$ and $\mu$ is a $(\delta,s,\delta^{-\epsilon})$-measure on $[0,1]^{d}$, then for all $s\in[0,d],$ we have
@@
For a nonnegative number $\sigma,$ a function $f:[0,1]\rightarrow\mathbb{R}$ will be called $\sigma$-superlinear on $[a,b]\subseteq[0,1]$, if for all $x\in[a,b],$ we have that $f(x)\geq f(a)+\sigma\cdot(x-a).$ An interval $[a,b]\subseteq[0,1]$ is $\tau$-allowable if $\tau\leq b-a\leq a.$ Moreover, given $[a,b]\subseteq[0,1]$, we say that a collection of pairwise disjoint intervals $[a_{j},b_{j}]\subseteq[a,b]$ is $\tau$-allowable if each $[a_{j},b_{j}]$ is $\tau$-allowable. Now, given $[a,b]\subseteq[0,1]$, define
@@
where the supremum is taken over all $\tau$-allowable collections $\{[a_{j},b_{j}]\}$ of sub-intervals of $[a,b]$ and numbers $\sigma_{j}\in[0,d]$ such that $f$ is $\sigma_{j}$-superlinear on $[a_{j},b_{j}]$. Given $u\in[0,d]$, define $\mathcal{L}_{d,u}$ to be the set of all non-decreasing Lipschitz functions $f:[0,1]\to[0,d]$ with Lipschitz constant $\leq d$ that are piecewise linear, with $f(0)=0$ and $f(x)\geq ux$ for all $x\in[0,1]$. Putting all of this together, we define
@@
[…]
@@
[Variant of Lemma 5.18 in [20]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $[0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has $(s,K,1-c)$-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
@@
For some constant $C_{0},$ only depending on the size of the support of $\mu_{1}$ and $\mu_{2}$, we have that by Proposition 6.10 and the method of constructing $D_{s},$ the measure $\mu\circ V_{x}$ is $(t\to D_{s}(t)-\eta)$-adapted at any scales in the range $[0,(K/C_{0})^{-C^{\prime}\eta^{-1}}]$ for any $\eta$ that is small enough. The fact that $\Sigma_{\tau}(D_{s};t,d)$ is continuous as a function in $t$ follows by Lemma 6.8, so by applying Proposition 6.11, we obtain that if $\eta$ is small enough in terms of $u$ and $\zeta,$ then
@@ 6.2.2 Modifications to their argument
Using Lemma 6.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(t,K,1-c,0)$-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
@@
[…]
@@
7 Proof of Theorem 1.2
Since $H^{s}(A)>0$, by the Frostman lemma [12, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$. Let $X$ and $Y$ be i.i.d. random variables that both have distribution measure $\mu$.
Lemma 5.1 gives us
@@
Let $\mathcal{N}_{2^{-n}}(A+A)$ and $\mathcal{N}_{2^{-n}}(\phi(A,A))$ be the least number of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively. It is clear that
$\log\mathcal{N}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~{}~{}\log\mathcal{N}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
@@
$\max\left(\frac{\log\mathcal{N}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{N}_{2^{-n}}(\phi(A,A))}{n}\right)$
$\quad\geq\max\left(\frac{H_{n}(X+Y)}{n},\frac{H_{n}(\phi(X,Y))}{n}\right)\geq(1+\epsilon)s.$
When $n\to\infty$, one obtains
$\max(\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A)))\geq(1+\epsilon)s.$
This completes the proof.
Appendix A Appendix - Examples

### v2

@@ 4.3 The general forms
In the conclusions of Lemma 4.1 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C_{1}C_{2})$, $O_{\phi,s}(\log_{+}C_{1}C_{2})$ or $O_{s}(\log_{+}C)$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $1/2<s<1$; (2) the range of random variables are bounded away from 0 by some absolute constant. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@
Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$. Let $(X^{\prime},Y^{\prime})$ be i.i.d. with $(X,Y)$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
@@
To compare Theorems 2.30 and 5.1, one may take $W=(X,Y)$ in Theorem 2.30. The random variables $X$ and $Y$ are required to be independent in Theorem 5.1, but may be dependent in Theorem 2.30.
@@
$\Delta(\mu^{\prime},\mu^{\prime\prime})(B)=\int\mu^{\prime}(\{x\colon|x-y|\in B\})d\mu^{\prime\prime}(y)$
@@
Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $0<s\leq 1/2$ and $C\geq 1$. Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$. Then for any $\zeta>0$, we have that
@@
Let $0<s<1$ and $C\geq 1$. Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$. For any $0<\gamma<1/3$, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
@@
By construction, $\text{dist}(U_{1},U_{2})\geq 2^{-k}\gg(\gamma/C)^{1/s}$. And by Lemma A.2, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman. This completes the proof. ∎
@@
Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later. We apply Lemma 5.3 on the distribution measure of $X$, denoted by $\mu$. Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
@@
For simplicity, write $a=\mu(U_{1})$ and $b=\mu(U_{2})$, with $1/2-\gamma\leq a,b\leq 1/2$. It follows that
@@ 5.2.1 Material from Shmerkin and Wang
We now record some related concepts that we will need in the results below. Basically, given two measures, $\mu$ and $\nu,$ these definitions will quantify how much mass from $\nu$ can assign to a projection associated to a point in the support of $\mu.$ Specifically, for $x\in\mathbb{R}^{d},$ and a measure $\nu\in\mathcal{P}(\mathbb{R}^{d}),$ the pair $(x,\nu)$ is said to have $(t,K,c,\delta)$-thin tubes if there exists a Borel function $\phi_{x}:\mathbb{R}^{d}\to\mathbb{R}$ such that $0\leq\phi_{x}\leq 1$, $\int\phi_{x}d\nu\geq c$, and
@@
[…]
@@
[Variant of Proposition 5.11 in [24]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in [0,1],[0,1],then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
@@
$P_{v}(\mu_{2}\times\nu)(B(z,r))=\int\nu\left(B\left(\frac{z-ax}{b},\frac{r}{b}\right)\right)d\mu_{2}(x)\leq\int C\left(\frac{r}{b}\right)^{s}d\mu_{2}(x)\leq 2Cr^{s}$
@@
$P_{v}(\mu_{2}\times\nu)(B(z,r))=\int\mu_{2}\left(B\left(\frac{z-by}{a},\frac{r}{a}\right)\right)d\nu(y)\leq\int C\left(\frac{r}{a}\right)^{s}d\nu(y)\leq 2Cr^{s}$
@@
Next we introduce a particularly useful a relaxation of the notion of an $(s,C)$-Frostman measure. In a metric space $X,$ we will call a measure $\mu\in\mathcal{P}(X)$ a $(\delta,s,C)$-measure if $\mu(B_{r})\leq C\cdot r^{s}$ for all $r\in[\delta,1].$ Given $t\in[0,d]$, $s\in[0,k],$ a function $D:[0,d]\to[0,k],$ and $C\geq 1$, we say that a measure $\rho\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$ is $(D;C)$-adapted at scale $\delta$ if whenever $\epsilon>0$ and $\mu$ is a $(\delta,s,\delta^{-\epsilon})$-measure on $0,1]^{d}$, then for all $s\in[0,d],$ we have
@@
For a nonnegative number $\sigma,$ a function $f:[0,1]\rightarrow\mathbb{R}$ will be called $\sigma$-superlinear on $a,b]\subseteq[0,1]$, if for all $x\in[a,b],$ we have that $f(x)\geq f(a)+\sigma\cdot(x-a).$ An interval $a,b]\subseteq[0,1]$ is $\tau$-allowable if $\tau\leq b-a\leq a.$ Moreover, given $a,b]\subseteq[0,1]$, we say that a collection of pairwise disjoint intervals $a_{j},b_{j}]\subseteq[a,b]$ is $\tau$-allowable if each $a_{j},b_{j}]$ is $\tau$-allowable. Now, given $a,b]\subseteq[0,1]$, define
@@
where the supremum is taken over all $\tau$-allowable collections $\{[a_{j},b_{j}]\}$ of sub-intervals of $[a,b]$ and numbers $\sigma_{j}\in[0,d]$ such that $f$ is $\sigma_{j}$-superlinear on $a_{j},b_{j}]$. Given $u\in[0,d]$, define $\mathcal{L}_{d,u}$ to be the set of all non-decreasing Lipschitz functions $f:[0,1]\to[0,d]$ with Lipschitz constant $\leq d$ that are piecewise linear, with $f(0)=0$ and $f(x)\geq ux$ for all $x\in[0,1]$. Putting all of this together, we define
@@
$F_{y}(\mu_{X_{y}})\text{ is }(2^{-m},\Sigma_{\tau}(D;s,d)-O(\epsilon^{1/2}),2^{-\eta m/2})\text{-robust, provided that},$
@@
[Variant of Lemma 5.18 in [24]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
@@
For some constant $C_{0},$ only depending on the size of the support of $\mu_{1}$ and $\mu_{2}$, we have that by Proposition 5.10 and the method of constructing $D_{s},$ the measure $\mu\circ V_{x}$ is $(t\to D_{s}(t)-\eta)$-adapted at any scales in the range $0,(K/C_{0})^{-C^{\prime}\eta^{-1}}]$ for any $\eta$ that is small enough. The fact that $\Sigma_{\tau}(D_{s};t,d)$ is continuous as a function in $t$ follows by Lemma 5.8, so by applying Proposition 5.11, we obtain that if $\eta$ is small enough in terms of $u$ and $\zeta,$ then
@@ 5.2.2 Modifications to their argument
Using Lemma 5.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
@@
$(\rho_{y})_{X_{y}^{\prime}}\text{ is }(\varphi(2s)-\zeta/2,\delta_{0}^{-1})\text{-Frostman}$ (57)
@@
6 Proof of Theorems 1.5-1.7
Proof of Theorem 1.5.
Since $H^{s}(A)>0$, by the Frostman lemma [13, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$. Let $X$ and $Y$ be i.i.d. random variables that both have distribution measure $\mu$.
Theorem 1.2 gives us
@@
Recall that $\mathcal{E}_{2^{-n}}(A+A)$ and $\mathcal{E}_{2^{-n}}(\phi(A,A))$ are the least numbers of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively. It is clear that
$\log\mathcal{E}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~~\log\mathcal{E}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
@@
$\max\left\{\frac{\log\mathcal{E}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{E}_{2^{-n}}(\phi(A,A))}{n}\right\}\geq\frac{1}{n}\max\left\{H_{n}(X+Y),H_{n}(\phi(X,Y))\right\}\geq(1+\epsilon)s.$
When $n\to\infty$, we obtain
$\max\{\underline{\dim}_{B}(A+A),\,\overline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s,$
and
$\max\{\overline{\dim}_{B}(A+A),\,\underline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s.$
This completes the proof. ∎
Proof of Theorem 1.6.
We still choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ satisfying $A_{\eta\delta}+A_{\eta\delta}$ and $\phi(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+A)_{\delta/10}$ and $(\phi(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Since $A$ is $\delta$-separated, the Lebesgue measure of $A_{\eta\delta}$ is $\gg_{\phi}\delta^{1-s}$. Let $X$ be a random variable that has uniform distribution on $A_{\eta\delta}$. Combining the non-concentration condition, one deduces that
$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\delta^{s-1},\quad&\text{if }0<r\leq\delta,\\ Cr^{s},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $Cr^{s}$. So $X$ is $(s,O_{\phi}(C))$-Frostman.
Let $Y$ be a random variable that is i.i.d. with $X$. Now similar relations as in (58), (61) and (62) hold. As a replacement of (60), we apply Theorem 1.2 to get an $\epsilon^{\prime}=\epsilon^{\prime}(s,\phi)$ such that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\}\geq n(s+\epsilon^{\prime})+O_{\phi,s}(\log C)+O_{\phi,s}(1).$
for all $n\in\mathbb{N}$. Putting everything together, we arrive at the conclusion that
$\max\{\mathcal{E}_{\delta}(A+A),\,\mathcal{E}_{\delta}(\phi(A,A))\}\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A)$
for all $\delta>0$. ∎
Proof of Theorem 1.7.
Let $\epsilon=\epsilon(s)$ to be determined later.
For any sets $E\subseteq{\mathbb{R}}$, $F\subseteq{\mathbb{R}}^{2}$, and any positive number $\lambda$, denote by $E_{\lambda}$ and $F_{\lambda}$ the rectangular $\lambda$-neighbourhood of $E$ and $F$, respectively. More concretely, $E_{\lambda}=\bigcup\nolimits_{x\in E}(x-\lambda,x+\lambda)$ and $F_{\lambda}=\bigcup\nolimits_{(x,y)\in F}(x-\lambda,x+\lambda)\times(y-\lambda,y+\lambda)$.
Note that $f_{1}(x,y)=x+y$ and $f_{2}(x,y)=\phi(x,y)=\alpha x^{2}+\beta y^{2}$ are $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\alpha\cdot A_{\eta\delta}^{\wedge 2}+\beta\cdot A_{\eta\delta}^{\wedge 2}$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})_{\delta/10}$, respectively, for any $\delta>0$. Then
$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\geq\frac{1}{2}\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})&\geq\frac{1}{2}\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2}).\end{split}$ (58)
Moreover, since $A$ is $\delta$-separated, we have
$\mathcal{E}_{\delta}(A_{\eta\delta})=\mathcal{E}_{\delta}(A)=\#A=\lceil\delta^{-s}\rceil.$ (59)
Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$. Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, the $2$-dimensional Lebesgue measure of $G_{\eta\delta}$ is $\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}$.
Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Combining the non-concentration condition, one can verify that
[…]
for any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$. In all the cases, the right-hand side can be bounded by $C^{2}\delta^{-\epsilon}r_{1}^{s}r_{2}^{s}$. So $(X,Y)$ is jointly $(s,s,C^{2}\delta^{-\epsilon})$-Frostman.
By Theorem 1.4, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (60)
for all $n\in\mathbb{N}$. In particular, we take $n$ such that $2^{-n-1}<\delta\leq 2^{-n}$.
By Lemma 2.4(ii), it follows that
$\begin{split}&\log\#\,range_{+}(D_{n}(X))\geq H_{n}(X),\quad\log\#\,range_{+}(D_{n}(Y))\geq H_{n}(Y)\\ &\qquad\qquad\quad\log\#\,range_{+}(D_{n}(X+Y))\geq H_{n}(X+Y),\\ &\qquad\quad\log\#\,range_{+}(D_{n}(\alpha x^{2}+\beta Y^{2})\big{\}}\geq H_{n}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (61)
Moreover,
$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\geq\frac{1}{2}\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\frac{1}{2}\#\,range_{+}(D_{n}(X+Y)),\\ &\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\mathcal{E}_{2^{-n}}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\frac{1}{2}\#\,range_{+}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (62)
Putting (58)-(62) together, one arrives at
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}$
$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1))}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{s}(\epsilon))}.$ (63)
Now we can take $\epsilon=\epsilon(s)>0$ to be sufficiently small so that $\epsilon^{\prime}-O_{s}(\epsilon)\geq\epsilon$ on the right-hand side of (63). Then
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
for any $\delta>0$.
Finally, by (59) one has $\mathcal{E}_{\delta}(A)\ll\delta^{-s}$. This gives
$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
whenever $\delta$ is sufficiently small depending on $\phi,s$ and $C$. The proof is completed. ∎
7 Examples